OpenStudy (annie96):

4. Complete the statement with a variable expression. Dale has three dollars more than twice the amount of money that Leo has. If Leo has d dollars, then Dale has ___ dollars. (Points : 1) A. 2d+3 B. d/2+3 C. 2(3)+d D. d+2+3

OpenStudy (annie96):

I need help with this problem if you don't mine. 5. Use the formula for the area of a rectangle A = lw to answer the question. A rectangle has a length of n units. Its width is 8 units more than its length. What is the area of the rectangle? (Points : 1) 2 + 8n n(n – 8) n(n + 8) 2n + 8

OpenStudy (texaschic101):

2D + 3 ========= A = lw l = n w = n + 8 now we sub A = n(n + 8)